# VITAL: More Understandable Feature Visualization through Distribution Alignment and Relevant Information Flow

### v1

@@ 6 Conclusion
Feature visualization (FV) is crucial for understanding neural network encoding, yet existing methods often produce repetitive patterns, irrelevant features, or artifacts. We reframe FV by optimizing for feature distribution similarity between generated and real images. We further suggest incorporating relevance scores to accurately capture the information that is encoded in an intermediate neuron, helping to avoid irrelevant features. Through diverse metrics and two human user studies, we show that VITAL provides clearer, more interpretable visualizations for both humans and machines while better capturing the information encoded in a target neuron. We anticipate VITAL to aid safety-critical domains like medicine and complement Mechanistic Interpretability by assigning meaning to mechanistic circuits.
@@ A.1 Implementation Details
[…]
@@
[…]
@@ Appendix B Additional Results
In this section, we provide more quantitative results (Sec. A.3, with various setups for methods Fourier [28], MACO [9], DeepInversion [44], and PII [12]. Furthermore, we provide more qualitative results for class neurons ( Figs. 17, 15, 16, 18, 19, 20, 21 and 22), intermediate neurons (Figs. 24 and 25), and results on disentangles polysemantic neurons (Fig. 26). Additionally, we extend our results with the performance of intermediate neurons (see Sec. B.1), analysis of LRP on class neurons (see Sec. B.2), concept-level visualization (see Sec. B.3), analysis on predictions (see Sec. B.4), scalability across different architectures (see Sec. B.5), and of the failure cases (see Sec. B.6).
@@ B.3 Visualization of Concepts
As part of Mechanistic Interpretability, people are interested in finding concept-based explanations of model behavior. These concepts might be feature directions encoded through multiple neurons in a layer, which can be, for example, discovered by CRAFT [10]. In VITAL, we obtain these directions as well as the images that highly activate the concepts through CRAFT. To optimize for feature directions, we modify the initialization of relevances of target neurons in LRP to reflect the weights given by the feature direction. Specifically, for each image’s feature map at the penultimate layer, we compute the cosine similarity with the concept direction vector. Then, we obtain the pixel location of the highest cosine similarity to assign the direction vector as the initial relevance score and apply LRP as in intermediate neuron visualization. Through this modification, VITAL can give meaning to these feature directions. We provide several examples for concept visualization in Figs. 27, 28, 29 and 30.
@@
[…]
@@ B.4 Analysis on Predictions
We further investigate the predictions of images produced from different methods on ResNet50 and observed that other methods including MACO seem to produce irrelevant features that mislead the model, producing predictions unrelated to the original class. We give two examples of prediction including the second-highest class score in Sec. B.3.
@@ B.5 Scalability Across Different Architectures
We demonstrate the scalability of VITAL by evaluating its performance across various architectures and conducting qualitative assessments (see Figs. 17, 18, 19, 20, 21 and 22). Our results highlight the robustness of our approach, whereas other methods fail to achieve similar adaptability and consistency across different network designs. Furthermore, t-SNE projections in Fig. 14 reveal a similar trend in the embedding space. VITAL is the only method that reliably position generated features at the center of their respective clusters, capturing distinct characteristics that are recognizable.
@@ C.3 Effect of the Transparency Map
Table 5: Comparison of class-specific samplinf size for the reference images on ResNet50.
@@ C.5 Alternative Attribution Methods
We extend VITAL for class neuron, intermediate neuron and concept visualizations on ResNet50 by incorporating Guided Backpropagation as an alternative to LRP, providing additional insights into feature attributions and model interpretability. We provide qualitative results to compare Guided Backpropagation with LRP for class neurons in Fig. 11, intermediate neurons in Figs. 24 and 25, and concepts in Figs. 27, 28, 29 and 30.
@@ Appendix D Human Interpretability Study
Analysis and Results. In Figs. 37, 38 and 39, we provide a fine-grained analysis of each question across all sections of our user study complementing the results in the main paper. In particular, we provide class- or concept-specific score distribution for each method. We observe that, as before, our method performs favorably across all three tasks compared to other methods for each question. Furthermore, we see that VITAL yields consistently good results, showing better results in almost all cases across all study sections. There are specific classes, such as specific animals, or ”hamburger” and ”grand piano”, where our method yields much more interpretable visualizations.
@@
Figure 14: t-SNE projection of embedding. We show a low-dimensional tSNE embedding of the features at the penultimate layer for five dog breeds indicated by color across different architectures. Transparent circles are original training images and FVs are indicated by symbols: $\blacksquare$: VITAL, $\blacktriangle$: MACO.

### v2

@@ 6 Conclusion
Feature visualization is key to understanding neural representations, but existing methods often yield artifacts or irrelevant patterns. We propose a new approach that aligns the feature distributions of generated and real images, and incorporates relevance scores to better capture the true information encoded in intermediate neurons. Across multiple metrics and two user studies, VITAL produces clearer and more interpretable visualizations for both humans and models. We envision VITAL supporting safety-critical fields such as medicine and complementing Mechanistic Interpretability by attributing meaning to neural circuits.
@@ A.1 Implementation Details
$\text{IndexX}=\text{torch.sort}(f_{l}(x),\text{dim}=2)$
$\text{SortedY},-=\text{torch.sort}(f_{l}(y),\text{dim}=2)$
$\text{SortedY}=\text{torch.mean}(\text{SortedY},\text{dim}=0)$
$\text{InverseIndex}=\text{IndexX.argsort}(-1)$
$g_{l}(y)=\text{SortedY.gather}(-1,\text{InverseIndex})$
@@
Figure 8: Overview of VITAL framework. Our VITAL framework mainly consists of two main components. In part (a), we utilize a pretrained frozen model () to generate visualizations. For the given type of visualization-class neurons or intermediate neurons-we first select NN reference images from the ImageNet training dataset, then we optimize a randomly initialized image (Refer to caption) through our feature distribution matching approach, which is applied across multiple layers of the model. For feature distribution matching, we utilize the feature activations $A$ and feature relevance (arrows) $A\odotRR$ for class neurons and intermediate neurons, respectively. We perform the feature distribution matching using (b) our sort matching loss, where we compute the difference between the feature distributions of $z$ (from synthetic data) and $z^{\prime}$ (from reference data). We achieve this by first sorting the values and obtaining the sorted indices as $\pi$ and $\pi^{\prime}$. Considering that sorting is a discrete operation and we want to allow back-propagation to optimize $z$, we obtain a reverse mapping $\bar{\pi}(\cdot)$ by sorting the indices $\pi^{\prime}$ to re-index $z^{\prime}$ to $z^{r}$, which allows $z$ to be unchanged. Thus, we were able to match $z$ and $z^{r}$ through minimizing the MSE loss. The proposed SM loss can be used in a plug-and-play manner without introducing any parameters, as summarized in Alg. 1.
@@ Appendix B Additional Results
In this section, we provide more quantitative results (Tab. 2, with various setups for methods Fourier [30], MACO [11], DeepInversion [46], and PII [14]. Furthermore, we provide more qualitative results for class neurons ( Figs. 20, 18, 19, 21, 22, 23, 24 and 25), intermediate neurons (Figs. 27 and 28), and results on disentangles polysemantic neurons (Fig. 29). Additionally, we extend our results with the performance of intermediate neurons (see Sec. B.1), analysis of LRP on class neurons (see Sec. B.2), concept-level visualization (see Sec. B.3), analysis on predictions (see Sec. B.4), scalability across different architectures (see Sec. B.5), and of the failure cases (see Sec. B.6).
@@ B.3 Visualization of Concepts
As part of Mechanistic Interpretability, people are interested in finding concept-based explanations of model behavior. These concepts might be feature directions encoded through multiple neurons in a layer, which can be, for example, discovered by CRAFT [12]. In VITAL, we obtain these directions as well as the images that highly activate the concepts through CRAFT. To optimize for feature directions, we modify the initialization of relevances of target neurons in LRP to reflect the weights given by the feature direction. Specifically, for each image’s feature map at the penultimate layer, we compute the cosine similarity with the concept direction vector. Then, we obtain the pixel location of the highest cosine similarity to assign the direction vector as the initial relevance score and apply LRP as in intermediate neuron visualization. Through this modification, VITAL can give meaning to these feature directions. We provide several examples for concept visualization in Figs. 30, 31, 32 and 33.
@@
MACO Tiger Apiary Silky terrier Coral reef
@@ B.4 Analysis on Predictions
We further investigate the predictions of images produced from different methods on ResNet50 and observed that other methods including MACO seem to produce irrelevant features that mislead the model, producing predictions unrelated to the original class. We give two examples of prediction including the second-highest class score in Tab. 4.
@@ B.5 Scalability Across Different Architectures
We demonstrate the scalability of VITAL by evaluating its performance across various architectures and conducting qualitative assessments (see Figs. 20, 21, 22, 23, 24 and 25). Our results highlight the robustness of our approach, whereas other methods fail to achieve similar adaptability and consistency across different network designs. Furthermore, t-SNE projections in Fig. 16 reveal a similar trend in the embedding space. VITAL is the only method that reliably position generated features at the center of their respective clusters, capturing distinct characteristics that are recognizable.
@@ C.3 Effect of the Transparency Map
Table 5: Comparison of class-specific sampling size for the reference images on ResNet50.
Figure 13: The qualitative impact of corrupting the reference images on class neuron visualization.
Acc. FID Zero-Shot Prediction
Top1 ($\uparrow$) ($\downarrow$) Top1 ($\uparrow$) Top5 ($\uparrow$)
Test Set 99.27 65.93 61.63 89.30
Train Set 99.90 58.79 66.62 92.56
Corrupt-5 99.97 72.73 61.03 88.80
Corrupt-10 99.83 86.73 56.20 85.50
Corrupt-20 99.93 113.30 47.30 77.40
CIFAR-10 - - 85.47 99.09
Fourier 5.97 34.87 11.40 53.80
VITAL 100.00 0.55 78.30 98.90
Table 6: The quantitative impact of corrupting the reference images, test-set analysis and CIFAR-10 analysis on class neuron visualization with ResNet50.
@@ C.4 Effect of the Reference Images
Additionally, we extended our embedding analysis to further validate the reference set strategy. As represented in Fig. 17, we clustered all training samples from three ImageNet classes into subgroups and generated representative VITAL visualizations per cluster using 50 nearest neighbor images per cluster. The resulting t-SNE plots show that VITAL images cover diverse intra-class modes without collapsing to a single mode, confirming that our approach preserves both local feature fidelity and global semantic diversity within the same class.
To verify that VITAL is not dependent on the training data and the dataset that is being used, we also consider test set examples for the reference and CIFAR-10 dataset, confirming that it can handle data beyond the original training data and the ImageNet dataset (Tab. 6). Finally, we performed a corruption experiment, in which we corrupted the reference set by gradually adding 5, 10, and 20 images from outside the class (Fig. 13 and Tab. 6). As expected, image quality and metrics degrade progressively with increasing contamination; however, even with partial corruption, VITAL visualizations remain considerably more stable than prior FV methods.
@@ C.5 Alternative Attribution Methods
We extend VITAL for class neuron, intermediate neuron and concept visualizations on ResNet50 by incorporating Guided Backpropagation as an alternative to LRP, providing additional insights into feature attributions and model interpretability. We provide qualitative results to compare Guided Backpropagation with LRP for class neurons in Fig. 12, intermediate neurons in Figs. 27 and 28, and concepts in Figs. 30, 31, 32 and 33.
@@ Appendix D Human Interpretability Study
Analysis and Results. In Figs. 40, 41 and 42, we provide a fine-grained analysis of each question across all sections of our user study complementing the results in the main paper. In particular, we provide class- or concept-specific score distribution for each method. We observe that, as before, our method performs favorably across all three tasks compared to other methods for each question. Furthermore, we see that VITAL yields consistently good results, showing better results in almost all cases across all study sections. There are specific classes, such as specific animals, or ”hamburger” and ”grand piano”, where our method yields much more interpretable visualizations.
@@
Figure 16: t-SNE projection of embedding. We show a low-dimensional tSNE embedding of the features at the penultimate layer for five dog breeds indicated by color across different architectures. Transparent circles are original training images and FVs are indicated by symbols:
$\blacksquare$: VITAL, $\blacktriangle$: MACO.
Figure 17: VITAL visualizations generated from clustered ImageNet training samples (3 classes shown) using 50 nearest-neighbor reference images per cluster (10 clusters). $\blacksquare$: VITAL, $\blacktriangle$: cluster center.